# QE4PE: Word-level Quality Estimation for Human Post-Editing

### v1

@@ Abstract
Word-level quality estimation (QE) detects erroneous spans in machine translations, which can direct and facilitate human post-editing. While the accuracy of word-level QE systems has been assessed extensively, their usability and downstream influence on the speed, quality and editing choices of human post-editing remain understudied. Our QE4PE study investigates the impact of word-level QE on machine translation (MT) post-editing in a realistic setting involving 42 professional post-editors across two translation directions. We compare four error-span highlight modalities, including supervised and uncertainty-based word-level QE methods, for identifying potential errors in the outputs of a state-of-the-art neural MT model. Post-editing effort and productivity are estimated by behavioral logs, while quality improvements are assessed by word- and segment-level human annotation. We find that domain, language and editors’ speed are critical factors in determining highlights’ effectiveness, with modest differences between human-made and automated QE highlights underlining a gap between accuracy and usability in professional workflows.
@@ 1 Introduction
Recent years saw a steady increase in the quality of machine translation (MT) systems and their widespread adoption in professional translation workflows (Kocmi et al., 2024a). Still, human post-editing of MT outputs remains a fundamental step to ensure high-quality translations, particularly for challenging textual domains requiring native fluency and specialized terminology (Liu et al., 2024). Quality estimation (QE) techniques were introduced to reduce post-editing effort by automatically identifying problematic MT outputs without the need for human-written reference translations and were quickly integrated into industry platforms (Tamchyna, 2021). Segment-level QE models correlate well with human perception of quality (Freitag et al., 2024) and exceed the performance of reference-based metrics in specific settings (Rei et al., 2021; Amrhein et al., 2022, 2023). On the other hand, word-level QE methods for identifying error spans requiring revision have received less attention in the past due to their modest agreement with human annotations, despite their promise for more granular and interpretable quality assessment in line with modern MT practices (Zerva et al., 2024). In particular, while the accuracy of these approaches is regularly assessed in evaluation campaigns, research has rarely focused on assessing the impact of such techniques in realistic post-editing workflows, with notable exceptions suggesting limited benefits (Shenoy et al., 2021; Eo et al., 2022). This hinders current QE evaluation practices: by foregoing experimental evaluation with human editors, it is implicitly assumed that word-level QE will become helpful once sufficient accuracy is reached, without accounting for the additional challenges towards a successful integration of these methods in post-editing workflows.
@@
In this study dubbed QE4PE (Quality Estimation for Post Editing), we address this gap by conducting a large-scale study with 42 professional translators for the EnglishItalian and EnglishDutch directions to measure the impact of word-level QE on editing quality, productivity and usability. We aim for a realistic and reproducible setup, employing the high-quality open-source NLLB 3.3B MT model (NLLB Team et al., 2024) to translate challenging documents from biomedical and social media domains. We then conduct a controlled evaluation of post-editing with error spans using four highlight modalities: a supervised state-of-the-art QE model trained on human error annotations (XCOMET, Guerreiro et al., 2024a), an unsupervised method leveraging the uncertainty of the MT model during generation, oracle error spans obtained from the consensus of previous human post-editors, and a no highlight baseline. The human post-editing is done in a simple online interface built for this study, GroTE, supporting the real-time logging of granular editing data for a quantitative assessment of editing effort and productivity across highlight modalities. We also survey professionals using an online questionnaire to collect qualitative feedback about the usability and quality of the MT model, as well as the interface and error span highlights. Finally, a subset of the original MT outputs and their post-edited variants is annotated following the MQM and ESA protocols (Lommel et al., 2014; Kocmi et al., 2024b) to verify quality improvements after post-editing. See Figure 1 for an overview of the study. More generally, our work embraces recent calls for an evaluation of translation technologies that is centered on users’ experience (Guerberof-Arenas and Moorkens, 2023; Savoldi et al., 2025).
@@
We release all data, code and the GroTE editing interface to foster future studies on the usability of error span highlighting techniques for other word-level QE methods and translation directions.111https://github.com/gsarti/qe4pe
@@ 2 Related Work
Human post-editing of MT outputs is increasingly common in professional translator workflows, as it was shown to increase the productivity of translators while preserving translation quality across multiple domains (Liu et al., 2024). However, many factors were found to influence the variability of post-editing productivity across setups, including MT quality (Zouhar et al., 2021b), interface familiarity (Läubli et al., 2021), individual variability and source-target languages typological similarity (Sarti et al., 2022). Studies evaluating the post-editing process generally focus on productivity, i.e. number of processed words/characters per minute, and the temporal, technical and cognitive dimensions of post-editing effort, operationalized through behavioral metrics such as editing time, keystrokes and pauses (Krings, 2001; Sarti et al., 2022). We adopt these metrics for the QE4PE study and relate them to different highlight modalities.
@@
The field of quality estimation was initially concerned with MT model uncertainty (Blatz et al., 2004; Specia et al., 2009), but in time began focusing on predicting translation quality even without using references (Turchi et al., 2013, 2014; Kepler et al., 2019; Thompson and Post, 2020 inter alia). Advances in segment- and word-level QE research are regularly assessed in annual WMT campaigns (Fomicheva et al., 2021; Zerva et al., 2022, 2024; Blain et al., 2023), where the best-performing QE systems are usually Transformer-based language models trained on human quality judgments, such as the popular COMET model suite (Rei et al., 2020, 2021, 2022). The widespread adoption of the fine-grained Multidimensional Quality Metrics scale (MQM, Lommel et al., 2014) prompted a paradigm shift in MT evaluation, leading to new QE metrics predicting quality at various granularity levels (Kocmi and Federmann, 2023; Fernandes et al., 2023; Guerreiro et al., 2024a). Aside from supervised models, unsupervised methods exploiting model uncertainty and its internal mechanisms were proposed as efficient alternatives to identify potential error spans for MT (Fomicheva et al., 2020; Dale et al., 2023; Xu et al., 2023; Himmi et al., 2024, surveyed by Leiter et al. 2024) and other language generation tasks (Kuhn et al., 2023; Farquhar et al., 2024). In this work, we compare the downstream effectiveness of state-of-the-art supervised and unsupervised word-level QE metrics for post-editing settings.
Human Assessment of Quality Estimation
Automatic QE methods are widely used in the translation industry for triaging automatic translations (Tamchyna, 2021). While QE usage has been found helpful to increase the confidence and speed of human assessment (Mehandru et al., 2023; Zouhar et al., 2024b), an incautious usage of these techniques can lead to users being misled (Zouhar et al., 2021a). Interfaces supporting word-level error highlights were developed for studying MT post-editing (Coppers et al., 2018; Herbig et al., 2020) and code reviewing (Sun et al., 2022; Vasconcelos et al., 2024), with results suggesting that striking the right balance of user-provided information is fundamental to improve the editing experience and prevent cognitive overload. Most similar to our study, Shenoy et al. (2021) investigated the effect of synthetic word-level QE highlights for EnglishGerman post-editing on Wikipedia data, concluding that current word-level QE accuracy was still insufficient to produce tangible productivity benefits. In this work, we expand the scope of such evaluation by including two translation directions, two challenging real-world text domains and state-of-the-art MT and QE systems and methods.
@@ 3.1 Structure of the Study
The pretask allows the core translators (12 per language direction, see Section 3.4) to familiarize themselves with the GroTE interface and text highlights. Before starting, all translators complete a questionnaire to provide demographic and professional information about their profile (Table 7). In the pretask, all translators work in an identical setup, post-editing a small set of documents similar to those of the main task with supervised highlights. We rank the translators into three groups (faster, average, slower) based on their speed from editing logs. Individuals from each group are then assigned randomly to highlight modalities to ensure an equal representation of editing speeds in each highlight modality. This procedure is repeated independently for both translation directions to balance subjective variability in the main task.
@@
This task, conducted in the two weeks following the pretask, covers the majority of the collected data and is the main object of study for the analyses of Section 4. In the main task, 24 core translators work on the same texts using the GroTE interface, but each with their assigned highlight modality (three translators per modality in each translation direction, as in Figure 1). After the main task, translators complete a questionnaire on the quality and usability of the MT outputs, the interface and, where applicable, word highlights.
@@
After Main, core translators are asked to post-edit an additional small set of related documents with GroTE, but this time working all with the no highlight modality. This step lets us obtain baseline editing patterns for each translator to estimate individual speed and editing differences across highlight modalities without the confounder of interface proficiency of Pre.
@@
Finally, a subset of 148 main task segments is selected randomly for manual annotation by six additional translators per direction (see Section 3.4). For each segment, the original MT output and all its post-edited versions are annotated with MQM error spans, including minor/major error severity and MQM error categories (Lommel et al., 2014). Moreover, the annotator proposes corrections for each error span, ultimately providing a 0-100 quality score. We adopt this scoring system, which closely follows the ESA protocol (Kocmi et al., 2024b), following recent results showing its effectiveness and efficiency.222See Figure 6 for an overview of setup and guidelines.
In summary, for each translation direction we collect 3 full sets of oracle post-edits, 12 full sets of edits with behavioral logs for pre, main and post task data, 12 pre- and post-task questionnaire responses, and 13 subsets of main task post-edits/MT outputs annotated with MQM error spans, corrections and segment-level ESA ratings.
@@ 3.2 Highlight Modalities
We conduct our study on four highlight modalities across two severity levels (minor and major errors). Using multiple severity levels follows the current MT evaluation practices (Freitag et al., 2021, 2024), and previous results showing that users tend to prefer more informative word-level highlights (Shenoy et al., 2021; Vasconcelos et al., 2024).
@@
In this setting, word-level error spans are obtained using XCOMET-XXL (Guerreiro et al., 2024a), which is a multilingual Transformer encoder (Goyal et al., 2021) further trained for joint word- and sentence-level QE prediction. We select XCOMET-XXL in light of its broad adoption, open accessibility and state-of-the-art performance in QE across several translation directions (Zerva et al., 2024). For the severity levels, we use the default labels predicted by the model, ignoring confidence scores and mapping critical labels to the major level.
@@
In this modality, we exploit the access to the MT model producing the original translations to obtain uncertainty-based highlights. We assess the agreement of two unsupervised QE methods employing the MT model log-probabilities with human post-edits, following recent results showing these can act as significant predictors of human editing time, and hence of the presence of potential issues (Lim et al., 2024). Scores were collected for the EnglishItalian and EnglishDutch directions of QE4PE Oracle post-edits and DivEMT (Sarti et al., 2022) to identify the best-performing method.333Unsupervised evaluation results are shown in Table 13. Token-level highlights are aggregated at the word level to match the granularity of other modalities. Results show a strong performance for the variance of next-token log-probabilities for 10 steps of Monte Carlo Dropout (Gal and Ghahramani, 2016), even surpassing the agreement of supervised metrics across both datasets. We ultimately select this method, which was previously shown to provide an efficient alternative to model ensembling for estimating epistemic uncertainty at the sentence level (Fomicheva et al., 2020).
@@ 3.3 Data and MT model
We begin by selecting two translation directions, EnglishItalian and EnglishDutch. We intentionally focus on out-of-English translations as they are generally more challenging for modern MT models (Kocmi et al., 2023). We aim to identify documents that are manageable for professional translators without domain-specific expertise but still prove challenging for our MT model to ensure a sufficient amount of error spans across modalities.
We begin by translating 3672 multi-segment English documents from the WMT23 General and Biomedical MT shared tasks (Kocmi et al., 2023; Neves et al., 2023) and MT test suites to Dutch and Italian. Then, XCOMET-XXL is used to produce a first set of segment-level QE scores and word-level error spans for all segments. To make the study tractable, we further narrow down the selection of documents according to several heuristics to ensure a realistic editing experience and a balanced occurrence of error spans (see Appendix A for details and data statistics). This procedure yields 351 documents, from which we manually select a subset of 64 documents (413 segments, 8744 source words per post-editor) across two domains:
@@
While the presence of multiple domains in the same task can render our post-editing setup less realistic, we deem it essential to test the cross-domain validity of our findings.
@@
Before producing highlights, we manually introduce 13 critical errors in main task segments to assess post-editing thoroughness. Errors are produced, for example, by negating statements, inverting the polarity of adjectives, inverting numbers, and corrupting acronyms. We replicate the errors in both translation directions to enable direct comparison. Most of these errors were correctly identified across all three highlight modalities (examples in Table 4).

### v2

@@ Abstract
Word-level quality estimation (QE) methods aim to detect erroneous spans in machine translations, which can direct and facilitate human post-editing. While the accuracy of word-level QE systems has been assessed extensively, their usability and downstream influence on the speed, quality and editing choices of human post-editing remain understudied. In this study, we investigate the impact of word-level QE on machine translation (MT) post-editing in a realistic setting involving 42 professional post-editors across two translation directions. We compare four error-span highlight modalities, including supervised and uncertainty-based word-level QE methods, for identifying potential errors in the outputs of a state-of-the-art neural MT model. Post-editing effort and productivity are estimated from behavioral logs, while quality improvements are assessed by word- and segment-level human annotation. We find that domain, language and editors’ speed are critical factors in determining highlights’ effectiveness, with modest differences between human-made and automated QE highlights underlining a gap between accuracy and usability in professional workflows.
@@ 1 Introduction
Recent years have seen a steady increase in the quality of machine translation (MT) systems and their widespread adoption in professional translation workflows (Kocmi et al., 2024a). Still, human post-editing of MT outputs remains a fundamental step to ensure high-quality translations, particularly for challenging textual domains requiring native fluency and specialized terminology (Liu et al., 2024). Quality estimation (QE) techniques were introduced to reduce post-editing effort by automatically identifying problematic MT outputs without the need for human-written reference translations and were quickly integrated into industry platforms (Tamchyna, 2021). Segment-level QE models correlate well with human perception of quality (Freitag et al., 2024) and exceed the performance of reference-based metrics in specific settings (Rei et al., 2021; Amrhein et al., 2022, 2023). On the other hand, word-level QE methods for identifying error spans requiring revision have received less attention in the past due to their modest agreement with human annotations, despite their promise for more granular and interpretable quality assessment in line with modern MT practices (Zerva et al., 2024). In particular, while the accuracy of these approaches is regularly assessed in evaluation campaigns, research has rarely focused on assessing the impact of such techniques in realistic post-editing workflows, with notable exceptions suggesting limited benefits (Shenoy et al., 2021; Eo et al., 2022). This hinders current QE evaluation practices: by foregoing experimental evaluation with human editors, it is implicitly assumed that word-level QE will become helpful once sufficient accuracy is reached, without accounting for the additional challenges towards a successful integration of these methods in post-editing workflows.
@@
In this study, which we dub QE4PE (Quality Estimation for Post Editing), we address this gap by conducting a large-scale study with 42 professional translators for the EnglishItalian and EnglishDutch directions to measure the impact of word-level QE on editing quality, productivity and usability. We aim for a realistic and reproducible setup, employing the high-quality open-source NLLB 3.3B MT model (NLLB Team et al., 2024) to translate challenging documents from biomedical and social media domains. We then conduct a controlled evaluation of post-editing with error spans in four highlight modalities, i.e. using highlights derived from four word-level QE methods: a supervised state-of-the-art QE model trained on human error annotations (XCOMET, Guerreiro et al., 2024), an unsupervised method leveraging the uncertainty of the MT model during generation, oracle error spans obtained from the consensus of previous human post-editors, and a no highlight baseline. The human post-editing is performed using GroTE, a simple online interface we built to support the real-time logging of granular editing data, enabling a quantitative assessment of editing effort and productivity across highlight modalities. We also survey professionals using an online questionnaire to collect qualitative feedback about the usability and quality of the MT model, as well as the interface and error span highlights. Finally, a subset of the original MT outputs and their post-edited variants is annotated following the MQM and ESA protocols (Lommel et al., 2013; Kocmi et al., 2024b) to verify quality improvements after post-editing. See Figure˜1 for an overview of the study. Our work represents a step towards an evaluation of translation technologies that is centered on users’ experience (Guerberof-Arenas and Moorkens, 2023; Savoldi et al., 2025).
@@
We release all data, code and the GroTE editing interface to foster future studies on the usability of error span highlighting techniques for other word-level QE methods and translation directions.111Data: hf.co/gsarti/qe4pe. Code: gsarti/qe4pe.
@@ 2 Related Work
Human post-editing of MT outputs is increasingly common in professional translator workflows, as it was shown to increase the productivity of translators while preserving translation quality across multiple domains (Liu et al., 2024). However, many factors were found to influence the variability of post-editing productivity across setups, including MT quality (Zouhar et al., 2021b), interface familiarity (Läubli et al., 2022), individual variability and source-target languages typological similarity (Sarti et al., 2022). Studies evaluating the post-editing process generally focus on productivity, i.e. number of processed words/characters per minute, and the temporal, technical and cognitive dimensions of post-editing effort, operationalized through behavioral metrics such as editing time, keystrokes and pauses (Krings, 2001; Sarti et al., 2022). We adopt these metrics for the QE4PE study and relate them to different highlight modalities.
@@
The field of quality estimation was initially concerned with MT model uncertainty (Blatz et al., 2004; Specia et al., 2009), but in time began focusing on predicting translation quality even without using references (Turchi et al., 2013, 2014; Kepler et al., 2019; Thompson and Post, 2020 inter alia). Advances in segment- and word-level QE research are regularly assessed in annual WMT campaigns (Fomicheva et al., 2021; Zerva et al., 2022, 2024; Blain et al., 2023), where the best-performing QE systems are usually Transformer-based language models trained on human quality judgments, such as the popular COMET model suite (Rei et al., 2020, 2021, 2022). The widespread adoption of the fine-grained Multidimensional Quality Metrics scale (MQM, Lommel et al., 2013) prompted a paradigm shift in MT evaluation, leading to new QE metrics predicting quality at various granularity levels (Kocmi and Federmann, 2023; Fernandes et al., 2023; Guerreiro et al., 2024). Aside from supervised models, unsupervised methods exploiting model uncertainty and its internal mechanisms were proposed as efficient alternatives to identify potential error spans in MT outputs (Fomicheva et al., 2020; Dale et al., 2023; Xu et al., 2023; Himmi et al., 2024, surveyed by Leiter et al. 2024). In this work, we compare the downstream effectiveness of state-of-the-art supervised and unsupervised word-level QE metrics for post-editing settings.
QE for Human Post-Editing Workflows
Automatic QE methods are widely used in the translation industry for triaging automatic translations (Tamchyna, 2021). While QE usage has been found helpful to increase the confidence and speed of human assessment (Mehandru et al., 2023; Zouhar et al., 2025), an incautious usage of these techniques can lead to a misplaced over-reliance on model predictions (Zouhar et al., 2021a). Interfaces supporting word-level error highlights were developed for studying MT post-editing (Coppers et al., 2018; Herbig et al., 2020) and code reviewing (Sun et al., 2022; Vasconcelos et al., 2025), with results suggesting that striking the right balance of user-provided information is fundamental to improve the editing experience and prevent cognitive overload. Most similar to our study, Shenoy et al. (2021) investigated the effect of synthetic word-level QE highlights for EnglishGerman post-editing on Wikipedia data, concluding that word-level QE accuracy was at the time still insufficient to produce tangible productivity benefits in human editing workflows. In this work, we expand the scope of such evaluation by including two translation directions, two challenging real-world text domains and state-of-the-art MT and QE systems and methods.
@@ 3.1 Structure of the Study
The pretask allows the core translators (12 per language direction, see Section˜3.4) to familiarize themselves with the GroTE interface and text highlights. Before starting, all translators complete a questionnaire to provide demographic and professional information about their profile (Table˜10). In the pretask, all translators work in an identical setup, post-editing a small set of documents similar to those of the main task with supervised highlights. We assign core translators into three groups based on their speed from editing logs (4 translators per group for faster, average and slower groups in each direction). Individuals from each group are then assigned randomly to each highlight modality to ensure an equal representation of editing speeds, resulting in 1 faster, 1 average and 1 slower translator for each highlight modality. This procedure is repeated independently for both translation directions.
@@
This task, conducted in the two weeks following the pretask, covers the majority of the collected data and is the main object of study for the analyses of Section˜4. In the main task, 24 core translators work on the same texts using the GroTE interface, with three translators per modality in each translation direction, as shown in Figure˜1. After the main task, translators complete a questionnaire on the quality and usability of the MT outputs, the interface and, where applicable, word highlights.222We do not disclose the highlight modality to translators to avoid biasing their judgment in the evaluation.
@@
After Main, the 12 core translators per direction are asked to post-edit an additional small set of related documents with GroTE, but this time working all with the no highlight modality. This step lets us obtain baseline editing patterns for each translator to estimate individual speed and editing differences across highlight modalities without the confounder of interface proficiency accounted for in the Pre stage.
@@
Finally, a subset consisting of 148 main task segments is randomly selected for manual annotation by six new translators per direction (see Section˜3.4). For each segment, the original MT output and all its post-edited versions are annotated with MQM error spans, including minor/major error severity and a subset of MQM error categories including e.g. mistranslations, omissions and stylistic errors (Lommel et al., 2013).333See Figure 5 for an overview of setup and guidelines. Moreover, the annotator proposes corrections for each error span, ultimately providing a 0-100 quality score matching the common DA scoring adopted in multiple WMT campaigns. We adopt this scoring system, which closely adheres to the ESA evaluation protocol (Kocmi et al., 2024b), following recent results showing its effectiveness and efficiency for ranking MT system.
In summary, for each translation direction, we collect 3 full sets of oracle post-edits, 12 full sets of edits with behavioral logs for pre, main and post task data, and 13 subsets of main task data (12 post-edits, plus the original MT output) annotated with MQM error spans, corrections and segment-level ESA ratings. Moreover, we also collect 12 pre- and post-task questionnaire responses from core set translators to obtain a qualitative view of the editing process.
@@ 3.2 Highlight Modalities
We conduct our study on four highlight modalities across two severity levels (minor and major errors). Using multiple severity levels follows the current MT evaluation practices (Freitag et al., 2021, 2024), and previous results showing that users tend to prefer more granular and informative word-level highlights (Shenoy et al., 2021; Vasconcelos et al., 2025). The highlight modalities we employ are:
@@
In this setting, word-level error spans are obtained using XCOMET-XXL (Guerreiro et al., 2024), which is a multilingual Transformer encoder (Goyal et al., 2021) further trained for joint word- and sentence-level QE prediction. We select XCOMET-XXL in light of its broad adoption, open accessibility and state-of-the-art performance in QE across several translation directions (Zerva et al., 2024). For the severity levels, we use the labels predicted by the model, mapping critical labels to the major level.
| Method | DivEMT | QE4PE | ||
| en-it | en-nl | en-it | en-nl | |
| Logprobs | 0.18 | 0.19 | 0.10 | 0.09 |
| MCD Var. | 0.41 | 0.42 | 0.23 | 0.31 |
| XCOMET (Sup.) | 0.34 | 0.35 | 0.16 | 0.19 |
| Avg. Trans. | - | - | 0.53 | 0.55 |
@@
In this modality, we exploit the access to the MT model producing the original translations to obtain uncertainty-based highlights. As a preliminary evaluation to select a capable unsupervised word-level QE method, we evaluate two unsupervised QE methods employing token log-probabilities assigned by MT model to predict human post-edits: raw negative log-probabilities (Logprobs), corresponding to the surprisal assigned by the MT model to every generated token, and their variance for 10 steps of Monte Carlo Dropout (MCD Var., Gal and Ghahramani, 2016). We employ surprisal-based metrics following previous work showing their effectiveness in predicting translation errors (Fomicheva and Specia, 2019) and human editing time (Lim et al., 2024). We collect scores for the EnglishItalian and EnglishDutch directions of QE4PE Oracle post-edits and DivEMT (Sarti et al., 2022) to identify the best-performing method, using metric scores extracted from the original models used for translation to predict human post-edits. We use average precision (AP) as a threshold-agnostic performance metric for the tested continuous methods. Oracle highlights obtained from the consensus of three annotator in the first stage of the study are used as reference for QE4PE, while a single set of post-edits is available for DivEMT. The XCOMET-XXL model used for Supervised highlights, and the average agreement of individual Oracle editors with the consensus label are also included for comparison. Table˜1444Full results in Table 16. Highlights are extended from tokens to words to match the granularity of other modalities. show a strong performance for the MCD Var. method, even surpassing the accuracy of the supervised XCOMET model across both datasets. Hence, we select MCD Var. for the Unsupervised highlight modality, setting value thresholds for minor/major errors to match the respective highlighted word proportions in the Supervised modality to ensure a fair comparison.
@@ 3.3 Data and MT model
| Task | Domain | # Docs | # Seg. | # Words |
| Pre | Social | 4 | 23 | 539 |
| Biomed. | 2 | 15 | 348 | |
| Main | Social | 30 | 160 | 3375 |
| Biomed. | 21 | 165 | 3384 | |
| Post | Social | 6 | 34 | 841 |
| Biomed. | 2 | 16 | 257 | |
| Total | 64 | 413 | 8744 | |
@@
We begin by selecting two translation directions, EnglishItalian and EnglishDutch, according to the availability of professional translators from our industrial partners. We intentionally focus on out-of-English translations as they are generally more challenging for modern MT models (Kocmi et al., 2023). We aim to identify documents that are manageable for professional translators without domain-specific expertise but still prove challenging for our MT model to ensure a sufficient amount of error spans across modalities. Since original references for our selected translation direction were not available, we do not have a direct mean to compare MT quality in the two languages. However, according to our human MQM assessment in Section˜4.3 (Table˜5), NLLB produces a comparable amount of errors across Dutch and Italian MT, suggesting similar quality.
We begin by translating 3,672 multi-segment English documents from the WMT23 General and Biomedical MT shared tasks (Kocmi et al., 2023; Neves et al., 2023) and MT test suites to Dutch and Italian. Our choice for these specialized domains, as opposed to e.g. generic news articles, is driven by the real-world needs of the translation industry for domain-specific post-editing support (Eschbach-Dymanus et al., 2024; Li et al., 2025). Moreover, focusing on domains that are considerably more challenging for MT systems than news, as shown by recent WMT campaigns (Neves et al., 2024), ensures a sufficient amount of MT errors to support a sound comparison of word-level QE methods. Then, XCOMET-XXL is used to produce a first set of segment-level QE scores and word-level error spans for all segments. To make the study tractable, we further narrow down the selection of documents according to several heuristics to ensure a realistic editing experience and a balanced occurrence of error spans (details in Appendix˜A). This procedure yields 351 documents, from which we manually select a subset of 64 documents (413 segments, 8,744 source words per post-editor) across two domains:
| Source en | So why is it that people jump through extra hoops to install Google Maps? |
| No High. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Oracle | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Sup. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| Unsup. | Quindi perché le persone devono fare un salto in più per installare Google Maps? |
| PENo High. | Quindi perché le persone devono fare un passaggio in più per installare Google Maps? |
| PEOracle | Allora, perché le persone fanno un passaggio in più per installare Google Maps? |
| PESup. | Quindi perché le persone fanno passaggi in più per installare Google Maps? |
| PEUnsup. | Quindi perché le persone fanno i salti mortali per installare Google Maps? |
@@
Table˜2 present statistics for the Pre, Main and Post editing stages, and Table˜3 shows an example of highlights and edits. While the presence of multiple domains in the same task can render our post-editing setup less realistic, we deem it essential to test the cross-domain validity of our findings.
@@
Before producing highlights, we manually introduce 13 critical errors in main task segments to assess post-editing thoroughness. Errors are produced, for example, by negating statements, inverting the polarity of adjectives, inverting numbers, and corrupting acronyms. We replicate the errors in both translation directions to enable direct comparison. Most of these errors were correctly identified across all three highlight modalities (examples in Table˜7).